OpenStudy (anonymous):

f(x)=sqrt(x^2+1)*cosx , find f'(x)

OpenStudy (anonymous):

\[f(x)= \sqrt{x^{2}+1}Cos(x)\] \[f\prime(x)=\frac{xCos(x)}{\sqrt{x^{2}+1}}-\sqrt{x^{2}+1}Sin(x) \;\;\;\;\; \] Use product rule....chaing rule is used for derivative of the square root stuff, where \[x^{2}+1\]. is the inside function more commonly known as "u".

OpenStudy (anonymous):

Thanks!